Problem 3.11 Show using a chain of logical equivalences that (p → r)A(q → r) pv q) →
Answer #1

Some formulas using here are

\sim means negation.

\equiv means equivalent to

  • a\rightarrow b\equiv \sim a\vee b
  • \left ( a\wedge b \right )\vee c\equiv \left ( a\vee c \right )\wedge \left ( b\vee c \right )
  • \sim \left ( a\vee b \right )\equiv \sim a\wedge \sim b

-------------------------------------------------------------------------------

\left ( p\rightarrow r \right )\wedge \left ( q\rightarrow r \right )\equiv \left ( \sim p\vee r \right )\wedge \left ( \sim q\vee r \right )

\equiv \left ( \sim p\wedge \sim q \right )\vee r

-~ (p V q) V r

\equiv \left ( p\vee q \right )\rightarrow r

Hence proved.

--------------------------------------------------------